# GUIDE: Guidance-based Incremental Learning with Diffusion Models

### 06 Mar 2024 v1

@@ 4.2 Guidance toward classes from current task
Our method builds upon the standard generative replay approach with a diffusion model. In each task $i$, we begin by training a class-conditional diffusion model $\epsilon_{\theta_{i}}(\mathbf{x}_{t},t,y)$ on currently available data $\mathcal{D}_{i}$ along with synthetic data samples from preceding tasks generated by the frozen previous diffusion model $\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y)$.
Subsequently, we train a classifier $f_{\phi_{i}}(y|\mathbf{x})$ on a combination of real data samples $\mathcal{D}_{i}$ and samples generated from the previous diffusion model $\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y)$. However, when generating each rehearsal sample for classifier training, we guide the diffusion sampling process to one of the classes from the current task $i$. Assuming that we want to generate a replayed sample from class $y_{i-1}$, we modify the diffusion denoising process at each time step $t$ according to the Eq. 6, as follows:
$\begin{split}{\hat{\epsilon}}_{\theta-1}(\mathbf{x}_{t},t,y_{i-1})&=\epsilon_{\theta-1}(\mathbf{x}_{t},t,y_{i-1})\\ &+s\nabla_{\mathbf{x}_{t}}{\ell}\left(f_{\phi_{i}}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t}),y_{i}\right).\end{split}$ (10)
@@
Intuitively, this modification steers the diffusion process towards examples from the current task as depicted in Fig. 1. Simultaneously, since we utilize only the previous frozen diffusion model that is not trained on classes from the current task, we consistently obtain samples from desired class $y_{i-1}$. Rehearsal examples obtained with the modified sampling process yield lower outputs for class $y_{i-1}$ in the current classifier $f_{\phi_{i}}(y|\mathbf{x})$ compared to the previous classifier $f_{\phi_{i-1}}(y|\mathbf{x})$. Hence, these examples can be interpreted as data samples that are more likely to be forgotten in continual training. The effect of our guidance technique resembles the Retrospective Adversarial Replay (RAR) (Kumari et al., 2022), where authors show that training the classifier on rehearsal samples similar to examples from the current task helps the model to learn the boundaries between tasks. This is also in line with observation by Toneva et al. (2018), who show that the sample’s distance from the decision border is correlated with the number of forgetting events.
@@ 5.1 Experimental setup
We evaluate our approach using CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) image datasets. We split the CIFAR-10 dataset in 2 or 5 equal tasks and the CIFAR-100 dataset in 5 or 10 equal tasks.
@@
For evaluation, we use two metrics commonly used in continual learning: average accuracy and average forgetting, defined as in (Chaudhry et al., 2018). For each task $i\in[1,\dots,T]$ we define average accuracy as:
$\bar{A}_{i}=\frac{1}{i}\sum_{j=1}^{i}A_{j}^{i},$ (12)
where $A_{j}^{i}$ denotes the accuracy of a model on $j$-th task after training on $i$ tasks. Analogously, average forgetting at task $i\in[1,\dots,T]$ can be defined as:
$\bar{F}_{i}=\begin{cases}0,&i=1\\ \frac{1}{i-1}\sum_{j=1}^{i-1}\max_{1\leq k\leq i}A_{j}^{k}-A_{j}^{i},&\text{otherwise}.\end{cases}$ (13)
Baseline methods
We compare our approach with state-of-the-art generative replay methods. For DGR with VAE (Shin et al., 2017), DGR with distillation (Van de Ven & Tolias, 2019), RTF (van de Ven & Tolias, 2018) and GFR (Liu et al., 2020a) we calculate the results on benchmarks that we use with the code provided by the authors. For BIR (Van de Ven et al., 2020), we freeze the encoder model after training on the first task. We also compare the results to the Joint training on all data and simple Fine-tuning with no rehearsal. As a soft upper bound of the proposed method, we present a Continual Joint setting, where we train the classifier continually with full access to all previous tasks (perfect rehearsal with infinite buffer size). In all methods using the diffusion model (including DDGR), we use the same number of DDIM sampling steps to obtain rehearsal samples. Importantly, we do not use any pre-training on external datasets.
5.2 Implementation details
Our training procedure is divided into two parts. First, in each task $i$, we train the class-conditional diffusion model $\epsilon_{\theta_{i}}(\mathbf{x}_{t},t,y)$ in a standard self-rehearsal approach, independently of the classifier. We use 1000 steps for generating the replay samples from $\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y)$. Then, we train the classifier $f_{\phi_{i}}(y|\mathbf{x})$, and in each task $i$, we load the diffusion model $\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y)$ already trained on the previous task. In the first task, we train the classifier solely on real data samples. New rehearsal samples are generated in each iteration of the classifier’s training. The mini-batches are balanced to ensure that each contains an equal number of samples from each class encountered so far. To generate replay samples from the diffusion model for classifier training, we use DDIM with 50 steps for the CIFAR-10 dataset and DDIM with 100 steps for the CIFAR-100 dataset. The training details and hyperparameters are described in Appendix C and in the code repository [^1] https://github.com/cywinski/guide.
5.3 Experimental results
We evaluate our method on a set of CL benchmarks with a comparative analysis conducted in relation to other generative rehearsal techniques. As presented in Tab. 1, our method outperforms other evaluated methods regarding the average accuracy after the last task $\bar{A}_{T}$ by a considerable margin on all benchmarks. Specifically, GUIDE notably improves upon the standard DGR with diffusion model on both average incremental accuracy and forgetting. It also outperforms DDGR, another GR approach that uses the diffusion model, proving the superiority of our sampling technique.
We further compare our method to feature replay methods (BIR and GFR). For BIR, we freeze the feature extractor model after the first task, while GFR significantly limits its updates to the feature distillation loss. In the case of the GFR, we observe less forgetting on the CIFAR-100 dataset, which is the effect of drastically limited plasticity. For a more detailed analysis, in Fig. 4 we thoroughly benchmark our method against the standard GR scenario and GFR method, presenting the accuracy on each encountered task after each training phase on CIFAR-100 with 5 tasks. Our approach significantly improved upon the standard GR scenario in terms of knowledge retention from preceding tasks. It indicates that training on rehearsal examples generated by GUIDE successfully mitigates forgetting. The decreased accuracy on the most recent task in our method can be interpreted through the lens of the stability-plasticity trade-off (Grossberg, 1982), highlighting that while our approach substantially reduces the forgetting of classifier, it does so at the expense of its ability to assimilate new information. Additionally, we observe that thanks to freezing the generative model, GFR can maintain the performance on the initial task notably better than our method. However, limited training of feature extractors on subsequent tasks leads to a substantial final performance drop.
@@ 5.3 Experimental results
Method CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
$T=2$ $T=5$ $T=5$ $T=10$ $T=2$ $T=5$ $T=5$ $T=10$
Joint 93.14 $\pm$ 0.16 72.32 $\pm$ 0.24 - - - -
Continual Joint 85.63 $\pm$ 0.39 86.41 $\pm$ 0.32 73.07 $\pm$ 0.01 64.15 $\pm$ 0.98 07.91 $\pm$ 0.67 02.90 $\pm$ 0.08 07.80 $\pm$ 0.55 06.67 $\pm$ 0.36
Fine-tuning 47.22 $\pm$ 0.06 18.95 $\pm$ 0.20 16.92 $\pm$ 0.03 9.12 $\pm$ 0.04 92.69 $\pm$ 0.06 94.65 $\pm$ 0.17 80.75 $\pm$ 0.22 87.67 $\pm$ 0.07
DGR VAE 60.24 $\pm$ 1.53 28.23 $\pm$ 3.84 19.66 $\pm$ 0.27 10.04 $\pm$ 0.17 43.91 $\pm$ 5.40 57.21 $\pm$ 9.82 42.10 $\pm$ 1.40 60.31 $\pm$ 4.80
DGR+distill 52.40 $\pm$ 2.58 27.83 $\pm$ 1.20 21.38 $\pm$ 0.61 13.94 $\pm$ 0.13 70.84 $\pm$ 6.35 43.43 $\pm$ 2.60 29.30 $\pm$ 0.40 21.15 $\pm$ 1.30
RTF 51.80 $\pm$ 2.56 30.36 $\pm$ 1.40 17.45 $\pm$ 0.28 12.80 $\pm$ 0.78 60.49 $\pm$ 5.54 51.77 $\pm$ 1.00 47.68 $\pm$ 0.80 45.21 $\pm$ 5.80
BIR 53.97 $\pm$ 0.97 36.41 $\pm$ 0.82 21.75 $\pm$ 0.08 15.26 $\pm$ 0.49 64.97 $\pm$ 2.15 65.28 $\pm$ 1.27 48.38 $\pm$ 0.44 53.08 $\pm$ 0.75
GFR 64.13 $\pm$ 0.88 26.70 $\pm$ 1.90 34.80 $\pm$ 0.26 21.90 $\pm$ 0.14 25.37 $\pm$ 6.62 49.29 $\pm$ 6.03 19.16 $\pm$ 0.55 17.44 $\pm$ 2.20
DDGR 80.03 $\pm$ 0.65 43.69 $\pm$ 2.60 28.11 $\pm$ 2.58 15.99 $\pm$ 1.08 22.45 $\pm$ 1.13 62.51 $\pm$ 3.84 60.62 $\pm$ 2.13 74.70 $\pm$ 1.79
DGR diffusion 77.43 $\pm$ 0.60 56.61 $\pm$ 1.85 28.25 $\pm$ 0.22 15.90 $\pm$ 1.01 26.32 $\pm$ 0.90 43.79 $\pm$ 0.40 68.70 $\pm$ 0.65 80.38 $\pm$ 1.34
GUIDE 81.29 $\pm$ 0.75 63.12 $\pm$ 0.45 41.66 $\pm$ 0.40 26.13 $\pm$ 0.29 14.79 $\pm$ 0.36 29.25 $\pm$ 1.15 44.30 $\pm$ 1.10 60.54 $\pm$ 0.82
@@ 5.4 Analysis of the proposed guidance
To demonstrate that our method produces samples near the decision boundary of a classifier, we propose to launch a simple adversarial attack on the generated samples in order to check how easy it is to change their class to the one from the current task. Concretely, we adapt the method introduced by Goodfellow et al. (2015), and modify each rehearsal sample that was generated during training $\hat{\mathbf{x}}$ as follows:
where $\epsilon=0.1$ and $y_{i}=\underset{c\in\mathcal{C}_{i}}{\mathrm{argmax}}\;f_{\phi_{i}}(y=c|\hat{\mathbf{x}})$. Then, we calculate the proportion of cases where the classifier’s prediction for the modified sample $\hat{\mathbf{x}}^{*}$ differs from its prediction for the original generated sample $\hat{\mathbf{x}}$. As shown in Tab. 2, we can change the classifier’s prediction much more frequently when we sample the replay examples according to our method. Since rehearsal examples generated with GUIDE are much more likely to be misclassified after a simple modification with predefined magnitude, this indicates that the modification of diffusion’s prediction introduced in our method successfully moves the generations closer to the classifier’s decision boundary.
@@ 6.1 The effect of changing classifier scale
In Fig. 5, we also observe that the scaling parameter introduces a trade-off between the stability and plasticity of the continually trained classifier. When we increase $s$, the accuracy on the previous task increases along with a drop in accuracy on the current task.
Figure 5: Classifier scale impact on forgetting and ability to acquire new information. Up to $s=0.2$, when we increase scale, we reduce the forgetting but also observe a drop in the accuracy on the second task $T_{2}$. When we use too large scale $s$, the quality of samples drops significantly, along with the accuracy on the previous task $T_{1}$. We further present this effect in Appendix A.
@@
Variant CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
$T=2$ $T=5$ $T=5$ $T=10$ $T=2$ $T=5$ $T=5$ $T=10$
DGR diffusion 77.43 $\pm$ 0.60 56.61 $\pm$ 1.85 28.25 $\pm$ 0.22 15.90 $\pm$ 1.01 26.32 $\pm$ 0.9 43.79 $\pm$ 0.4 68.70 $\pm$ 0.65 80.38 $\pm$ 1.34
PREV + 80.03 $\pm$ 0.65 60.31 $\pm$ 0.44 31.35 $\pm$ 0.66 18.22 $\pm$ 0.52 22.45 $\pm$ 1.13 40.00 $\pm$ 0.6 64.80 $\pm$ 1.00 77.60 $\pm$ 0.50
PREV - 75.79 $\pm$ 1.20 57.05 $\pm$ 0.43 28.40 $\pm$ 0.04 15.79 $\pm$ 0.21 27.58 $\pm$ 0.95 44.8 $\pm$ 0.9 68.34 $\pm$ 0.29 80.42 $\pm$ 0.44
CURR - 78.72 $\pm$ 0.58 57.72 $\pm$ 0.95 30.57 $\pm$ 0.33 16.87 $\pm$ 0.83 23.89 $\pm$ 0.44 43.31 $\pm$ 1.5 65.42 $\pm$ 0.81 78.96 $\pm$ 1.16
GUIDE 81.29 $\pm$ 0.75 63.12 $\pm$ 0.45 41.66 $\pm$ 0.40 26.13 $\pm$ 0.29 14.79 $\pm$ 0.36 29.25 $\pm$ 1.15 44.30 $\pm$ 1.10 60.54 $\pm$ 0.82
6.2 Alternative variants of guidance
Guidance toward classes from previous tasks
@@ 6.2 Alternative variants of guidance
$\begin{split}{\hat{\epsilon}}_{\theta-1}(\mathbf{x}_{t},t,y_{i-1})&=\epsilon_{\theta-1}(\mathbf{x}_{t},t,y_{i-1})\\ &+s\nabla_{\mathbf{x}_{t}}{\ell}\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}f_{\phi_{i-1}}}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y_{i-1}}\right),\end{split}$ (15)
where $y_{i-1}$ denotes the class label from one of the previous tasks. As noted by (Dhariwal & Nichol, 2021), the application of classifier guidance creates a trade-off: it enhances the quality of the generated samples at the cost of their diversity. This approach is similar to the one introduced by Gao & Liu (2023) except that we do not use guidance in the process of continual diffusion training but only in the classifier’s training. Intuitively similar buffer-based methods are based on the herding algorithm (Welling, 2009) and used in iCaRL method (Rebuffi et al., 2017), which seeks to store samples that best represent the mean of classes in the feature space.
@@
Alternatively, we can guide the diffusion-denoising process in the opposite direction by maximizing the entropy of the classifier instead of minimizing it. As noted by Sehwag et al. (2022), such an approach steers the denoising diffusion process away from the high-density regions of the data manifold. Consequently, it should generate synthetic samples that resemble the rare instances in the training dataset, which are typically more challenging for the classifier to identify. In the first variant, we propose to use the previous classifier $f_{\phi_{i-1}}(y|\mathbf{x})$ to guide away from the old classes:
$\begin{split}{\hat{\epsilon}}_{\theta-1}(\mathbf{x}_{t},t,y_{i-1})&=\epsilon_{\theta-1}(\mathbf{x}_{t},t,y_{i-1})\\ &{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}-}s\nabla_{\mathbf{x}_{t}}{\ell}\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}f_{\phi_{i-1}}}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y_{i-1}}\right).\end{split}$ (16)

### 31 May 2024 v2

@@ 4.2 Guidance toward classes from current task
Our method builds upon the standard generative replay approach with a diffusion model. In each task $i$, we begin by training a classifier $f_{\phi_{i}}(y|\mathbf{x})$ on a combination of real data samples $\mathcal{D}_{i}$ and samples generated from the frozen previous diffusion model $\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y)$. Subsequently, we train a class-conditional diffusion model $\epsilon_{\theta_{i}}(\mathbf{x}_{t},t,y)$ on currently available data $\mathcal{D}_{i}$ along with synthetic data samples from preceding tasks generated by the previous diffusion model $\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y)$.
When generating each rehearsal sample for classifier training, we guide the diffusion sampling process to one of the classes from the current task $i$.
Assuming that we want to generate a replayed sample from class $y_{i-1}$, we modify the diffusion denoising process at each time step $t$ according to the Eq. 5, as follows:
${\hat{\epsilon}}_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})=\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})+s\nabla_{\mathbf{x}_{t}}{\ell}\left(f_{\phi_{i}}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t}),y_{i}\right).$ (9)
@@
Intuitively, this modification steers the diffusion process towards examples from the current task, as depicted in Fig. 1. Simultaneously, since we utilize only the previous frozen diffusion model that is not trained on classes from the current task, we consistently obtain samples from desired class $y_{i-1}$. Rehearsal examples obtained with the modified sampling process yield lower outputs for class $y_{i-1}$ in the current classifier $f_{\phi_{i}}(y|\mathbf{x})$ compared to the previous classifier $f_{\phi_{i-1}}(y|\mathbf{x})$. Hence, these examples can be interpreted as data samples that are more likely to be forgotten during continual training. We experimentally validate this statement in Sec. 5.4. The effect of our guidance technique resembles the idea introduced in Retrospective Adversarial Replay (RAR) [26], where authors show that training the classifier on rehearsal samples similar to examples from the current task helps the model to learn the boundaries between tasks. This is also in line with observation by Toneva et al. [44], who show that the sample’s distance from the decision border is correlated with the number of forgetting events.
@@ 5.1 Experimental setup
We evaluate our approach on CIFAR-10 and CIFAR-100 [25] image datasets. We split the CIFAR-10 dataset into 2 and 5 equal tasks and the CIFAR-100 dataset into 5 and 10 equal tasks. Moreover, to validate if our method can be extended to datasets with higher resolution, we also evaluate it on the ImageNet-100 [6] dataset split into 5 tasks.
@@
For evaluation on each task $i\in[1,\dots,T]$, we use two metrics commonly used in continual learning: average accuracy $\bar{A}_{i}=\frac{1}{i}\sum_{j=1}^{i}A_{j}^{i}$ and average forgetting $\bar{F}_{i}=\frac{1}{i-1}\sum_{j=1}^{i-1}\max_{1\leq k\leq i}(A_{j}^{k}-A_{j}^{i})$, where $A_{j}^{i}$ denotes the accuracy of a model on $j$-th task after training on $i$ tasks. We follow the definitions of both metrics from [5].
@@
Method CIFAR-10 CIFAR-100 ImageNet100 CIFAR-10 CIFAR-100 ImageNet100
$T=2$ $T=5$ $T=5$ $T=10$ $T=5$ $T=2$ $T=5$ $T=5$ $T=10$ $T=5$
Joint 93.14 $\pm$ 0.16 72.32 $\pm$ 0.24 66.85 $\pm$ 2.25 - - - - -
Continual Joint 85.63 $\pm$ 0.39 86.41 $\pm$ 0.32 73.07 $\pm$ 0.01 64.15 $\pm$ 0.98 50.59 $\pm$ 0.35 07.91 $\pm$ 0.67 02.90 $\pm$ 0.08 07.80 $\pm$ 0.55 06.67 $\pm$ 0.36 12.28 $\pm$ 0.07
Fine-tuning 47.22 $\pm$ 0.06 18.95 $\pm$ 0.20 16.92 $\pm$ 0.03 9.12 $\pm$ 0.04 13.49 $\pm$ 0.18 92.69 $\pm$ 0.06 94.65 $\pm$ 0.17 80.75 $\pm$ 0.22 87.67 $\pm$ 0.07 64.93 $\pm$ 0.00
DGR VAE 60.24 $\pm$ 1.53 28.23 $\pm$ 3.84 19.66 $\pm$ 0.27 10.04 $\pm$ 0.17 9.54 $\pm$ 0.26 43.91 $\pm$ 5.40 57.21 $\pm$ 9.82 42.10 $\pm$ 1.40 60.31 $\pm$ 4.80 40.26 $\pm$ 0.91
DGR+distill 52.40 $\pm$ 2.58 27.83 $\pm$ 1.20 21.38 $\pm$ 0.61 13.94 $\pm$ 0.13 11.77 $\pm$ 0.47 70.84 $\pm$ 6.35 43.43 $\pm$ 2.60 29.30 $\pm$ 0.40 21.15 $\pm$ 1.30 41.17 $\pm$ 0.43
RTF 51.80 $\pm$ 2.56 30.36 $\pm$ 1.40 17.45 $\pm$ 0.28 12.80 $\pm$ 0.78 8.03 $\pm$ 0.05 60.49 $\pm$ 5.54 51.77 $\pm$ 1.00 47.68 $\pm$ 0.80 45.21 $\pm$ 5.80 41.2 $\pm$ 0.20
MeRGAN 50.54 $\pm$ 0.08 51.65 $\pm$ 0.40 09.65 $\pm$ 0.14 12.34 $\pm$ 0.15 - - - - - -
BIR 53.97 $\pm$ 0.97 36.41 $\pm$ 0.82 21.75 $\pm$ 0.08 15.26 $\pm$ 0.49 8.63 $\pm$ 0.19 64.97 $\pm$ 2.15 65.28 $\pm$ 1.27 48.38 $\pm$ 0.44 53.08 $\pm$ 0.75 40.99 $\pm$ 0.36
GFR 64.13 $\pm$ 0.88 26.70 $\pm$ 1.90 34.80 $\pm$ 0.26 21.90 $\pm$ 0.14 32.95 $\pm$ 0.35 25.37 $\pm$ 6.62 49.29 $\pm$ 6.03 19.16 $\pm$ 0.55 17.44 $\pm$ 2.20 20.37 $\pm$ 1.47
DDGR 80.03 $\pm$ 0.65 43.69 $\pm$ 2.60 28.11 $\pm$ 2.58 15.99 $\pm$ 1.08 25.59 $\pm$ 2.29 22.45 $\pm$ 1.13 62.51 $\pm$ 3.84 60.62 $\pm$ 2.13 74.70 $\pm$ 1.79 49.52 $\pm$ 2.52
DGR diffusion 77.43 $\pm$ 0.60 59.00 $\pm$ 0.57 28.25 $\pm$ 0.22 15.90 $\pm$ 1.01 23.92 $\pm$ 0.92 26.32 $\pm$ 0.90 40.38 $\pm$ 0.32 68.70 $\pm$ 0.65 80.38 $\pm$ 1.34 54.44 $\pm$ 0.14
GUIDE 81.29 $\pm$ 0.75 64.47 $\pm$ 0.45 41.66 $\pm$ 0.40 26.13 $\pm$ 0.29 39.07 $\pm$ 1.37 14.79 $\pm$ 0.36 24.84 $\pm$ 0.05 44.30 $\pm$ 1.10 60.54 $\pm$ 0.82 27.60 $\pm$ 3.28
Baseline methods
We compare our approach with state-of-the-art generative replay methods. For fairness, in the evaluation of BIR [46], we freeze the encoder model after training on the first task. In order to evaluate MeRGAN [50] on generative replay scenario, we generate samples from the final generator network trained sequentially on all tasks to construct a training dataset for the ResNet18 classifier. We also compare the results to the Joint training on all data and simple Fine-tuning with no rehearsal. As a soft upper bound of the proposed method, we present a Continual Joint setting, where we train the classifier continually with full access to all previous tasks (perfect rehearsal with infinite buffer size). In all methods using the diffusion model (including DDGR), we use the same number of denoising steps to obtain rehearsal samples. We recalculated scores for all related methods using the code provided by the authors. Importantly, we do not use any pre-training on external datasets.
5.2 Implementation details
Our training procedure is divided into two parts. First, in each task $i$, we train the classifier $f_{\phi_{i}}(y|\mathbf{x})$. In order to generate rehearsal samples, we load the diffusion model $\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y)$ already trained on the previous task. Then, we train a class-conditional diffusion model $\epsilon_{\theta_{i}}(\mathbf{x}_{t},t,y)$ in a standard self-rehearsal approach, independently of the classifier. In the first task, we train both the diffusion model and the classifier solely on real data samples.
In our method, it is essential to generate rehearsal examples progressively as the classifier is trained on the task. This way, similarly to active learning techniques, at each step, we guide the generation process toward the most challenging samples – those that are close to the decision boundary of the classifier at the given moment. The mini-batches are balanced to ensure that each contains an equal number of samples from each class encountered so far. For clarity, we present a detailed pseudocode of the entire training procedure in Appendix C. All training hyperparameters can be found in Appendix D and in the code repository [^1] https://github.com/cywinski/guide.
5.3 Experimental results
@@
We evaluate our method on a set of CL benchmarks with a comparative analysis conducted in relation to other generative rehearsal techniques. In Tab. 1, we present the mean and standard deviation of results calculated for 3 random seeds. Our method outperforms other evaluated methods regarding the average accuracy after the last task $\bar{A}_{T}$ by a considerable margin on all benchmarks. Specifically, GUIDE notably improves upon the standard DGR with diffusion model on both average incremental accuracy and forgetting. It also outperforms DDGR, another GR approach that uses the diffusion model, proving the superiority of our sampling technique.
We further compare our method to feature replay methods (BIR and GFR). In the case of the GFR, we observe less forgetting on the CIFAR-100 and ImageNet100 datasets. This is the effect of the drastically limited plasticity of GFR due to the decrease in the number of updates to the feature extractor after the first task. For a more detailed analysis, in Fig. 3 we thoroughly benchmark our method against the standard GR scenario and GFR method, presenting the accuracy on each encountered task after each training phase on CIFAR-100 with 5 tasks. Our approach significantly improved upon the standard GR scenario in terms of knowledge retention from preceding tasks. It indicates that training on rehearsal examples generated by GUIDE successfully mitigates forgetting. The decreased accuracy on the most recent task in our method can be interpreted through the lens of the stability-plasticity trade-off [15], highlighting that while our approach substantially reduces the forgetting of classifier, it does so at the expense of its ability to assimilate new information. Although GFR can maintain the performance on the initial task notably better than our method thanks to limited training of feature extractor on subsequent tasks, it leads to a substantial final performance drop.
@@ 5.4 Analysis of the proposed guidance
To demonstrate that our method produces samples near the decision boundary of a classifier, we propose to launch a simple adversarial attack on the generated samples in order to check how easy it is to change their class to the one from the current task. Concretely, we adapt the method introduced by Goodfellow et al. [14], and modify each rehearsal sample that was generated during training $\hat{\mathbf{x}}$ as follows:
where $\epsilon=0.1$ and $y_{i}=\underset{c\in\mathcal{C}_{i}}{\mathrm{argmax}}\;f_{\phi_{i}}(y=c|\hat{\mathbf{x}})$.
Then, we calculate the proportion of cases where the classifier’s prediction for the modified sample $\hat{\mathbf{x}}^{*}$ differs from its prediction for the original generated sample $\hat{\mathbf{x}}$. As shown in Tab. 2, we can change the classifier’s prediction much more frequently when we sample the replay examples according to our method. Since rehearsal examples generated with GUIDE are much more likely to be misclassified after a simple modification with predefined magnitude, this indicates that the modification of diffusion’s prediction introduced in our method successfully moves the generations closer to the classifier’s decision boundary.
@@ 6.1 The effect of changing classifier scale
In Fig. 5, we present results averaged over 3 random seeds for different values of gradient scale $s$, presenting standard deviation as error bars. We observe that the scaling parameter introduces a trade-off between the stability and plasticity of the continually trained classifier. When we increase $s$, the accuracy on the previous task increases along with a drop in accuracy on the current task.
6.2 Alternative variants of guidance
@@
Variant CIFAR-10 CIFAR-100 ImageNet100 CIFAR-10 CIFAR-100 ImageNet100
$T=2$ $T=5$ $T=5$ $T=10$ $T=5$ $T=2$ $T=5$ $T=5$ $T=10$ $T=5$
DGR diffusion 77.43 $\pm$ 0.60 59.00 $\pm$ 0.57 28.25 $\pm$ 0.22 15.90 $\pm$ 1.01 23.92 $\pm$ 0.92 26.32 $\pm$ 0.90 40.38 $\pm$ 0.32 68.70 $\pm$ 0.65 80.38 $\pm$ 1.34 54.44 $\pm$ 0.14
PREV + 80.03 $\pm$ 0.65 60.31 $\pm$ 0.44 31.35 $\pm$ 0.66 18.22 $\pm$ 0.52 29.13 $\pm$ 2.81 22.45 $\pm$ 1.13 40.00 $\pm$ 0.60 64.80 $\pm$ 1.00 77.60 $\pm$ 0.50 45.15 $\pm$ 3.64
PREV - 75.79 $\pm$ 1.20 57.05 $\pm$ 0.43 28.40 $\pm$ 0.04 15.79 $\pm$ 0.21 12.60 $\pm$ 0.71 27.58 $\pm$ 0.95 44.80 $\pm$ 0.90 68.34 $\pm$ 0.29 80.42 $\pm$ 0.44 59.98 $\pm$ 0.28
CURR - 78.72 $\pm$ 0.58 57.72 $\pm$ 0.95 30.57 $\pm$ 0.33 16.87 $\pm$ 0.83 23.55 $\pm$ 3.27 23.89 $\pm$ 0.44 43.31 $\pm$ 1.50 65.42 $\pm$ 0.81 78.96 $\pm$ 1.16 53.85 $\pm$ 3.60
GUIDE 81.29 $\pm$ 0.75 64.47 $\pm$ 0.45 41.66 $\pm$ 0.40 26.13 $\pm$ 0.29 39.07 $\pm$ 1.37 14.79 $\pm$ 0.36 24.84 $\pm$ 0.05 44.30 $\pm$ 1.10 60.54 $\pm$ 0.82 27.60 $\pm$ 3.28
Guidance towards classes from previous tasks
@@
${\hat{\epsilon}}_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})=\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})+s\nabla_{\mathbf{x}_{t}}{\ell}\left({\color[rgb]{0,0,1}f_{\phi_{i-1}}}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),{\color[rgb]{0,0,1}y_{i-1}}\right),$ (12)
where $y_{i-1}$ denotes the class label from one of the previous tasks. As noted by [8], the application of classifier guidance creates a trade-off: it enhances the quality of the generated samples at the cost of their diversity. This approach is similar to the one introduced by Gao and Liu [12] except that we do not use guidance in the process of continual diffusion training but only in the classifier’s training. Intuitively similar buffer-based methods are based on the herding algorithm [49] and used in iCaRL method [36], which seeks to store samples that best represent the mean of classes in the feature space.
@@
Alternatively, we can guide the diffusion-denoising process in the opposite direction by maximizing the entropy of the classifier instead of minimizing it. As noted by Sehwag et al. [40], such an approach steers the denoising diffusion process away from the high-density regions of the data manifold. Consequently, it should generate synthetic samples that resemble the rare instances in the training dataset, which are typically more challenging for the classifier to identify. In the first variant, we propose to use the previous classifier $f_{\phi_{i-1}}(y|\mathbf{x})$ to guide away from the old classes:
${\hat{\epsilon}}_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})=\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1}){\color[rgb]{0,0,1}-}s\nabla_{\mathbf{x}_{t}}{\ell}\left({\color[rgb]{0,0,1}f_{\phi_{i-1}}}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),{\color[rgb]{0,0,1}y_{i-1}}\right).$ (13)